# Nonlinear Causality in Brain Networks: With Application to Motor Imagery vs Execution

### v1

@@ 2.1 Inference in high-dimensional networks of time series
Most phenomena in the real world are composed of interconnected components. Temporal observations of multi-component systems form multivariate time series datasets. In the context of this study, it is appropriate to refer to this data structure as a network of time series, as this terminology is more inclusive and clearly emphasizes interconnectedness. Statistical inference in such high-dimensional networks of time series continues to pose a significant analytical challenge. For instance, in an Economics study designed to assess the economic conditions of a country or countries, it will be essential to analyze datasets of time-dependent indicators. In Finance, the problem of identifying the best asset allocation among all possible stocks would require analyzing a complex network of time series. Or, for example, to study climate change, it would be necessary to work on a network of time series of different meteorological variables from many spatial locations. Ombao and Pinto (2022) equip a unified framework with statistical methods for addressing such networks of time series based problems.
In such interconnected multi-node networks, if the primary purpose and scope of the research are to determine the connectivity between any specific two nodes, the confounding effects of the other nodes in the network cannot be ignored. In this case, a multi-parameter model to analyze the whole network or a method to eliminate the effects of other nodes in the network should be followed. Since the estimation and interpretation of multi-parameter models become more difficult as the number of parameters increases, instead of using a multi-parameter model, we adopt a practical and innovative dimensionality reduction approach proposed by Aslan and Ombao (2024) to determine the connectivity between two specific nodes. This approach employs spectral dynamic principal component analysis (sDPCA) in the frequency domain to create a low-dimensional representation of all other nodes. The resulting principal component scores are subsequently removed from the nodes of interest, thus eliminating the confounding effects. This method maintains the temporal and spectral characteristics of the data, which is crucial for accurate connectivity analysis. It ensures that the effects of other nodes are effectively isolated, allowing for a precise determination of connectivity between the two nodes.
Define $N=\{Y^{*}_{t},X^{*}_{t},Z_{1,t},\ldots,Z_{n-2,t}\}$ to be a network of time series with $n$ nodes. In this paper, we consider a network representing brain activity through EEG recordings, where each component (i.e., node or channel) is a time series. Within this network, denoted as $N$, $Y^{*}$, and $X^{*}$ represent any two nodes of interest. For problems that require component-wise solutions, our objective is to investigate the interaction or effective connectivity between any pair of nodes rather than the relationships within the entire network. These two nodes are influenced by other nodes due to the network’s interconnected structure. Therefore, to accurately analyze the effective connection between any two nodes of interest, it is necessary to isolate them from the influences of other nodes. As demonstrated and proposed by Aslan and Ombao (2024), sDPCA in the frequency domain offers practical benefits and is highly effective for the alleviation of dimensionality problems in high-dimensional settings. Thereby, the interference-free counterparts of the nodes of interest can be obtained in the following manner:
@@
where $f_{z}(.)$ is a function that accounts for the dynamic variability of the excluded nodes. In practice, $f_{z}(.)$ serves as a filter and can take any form, provided it effectively summarizes the net discharge of the excluded nodes. See Brillinger (1974); Ombao and Ho (2006); Hörmann et al. (2015); Hörmann and Kidzinski (2024); Wang et al. (2019); Ombao and Pinto (2022); Aslan and Ombao (2024) for more details.
@@ 2.2 Background on the threshold autoregressive model
The concept of switching between regimes based on a change point was first introduced by Quandt (1958) in linear regression. Two decades later, Tong (1978) and Tong and Lim (1980) extended this idea to nonlinear time-series analysis through threshold non-linearity. Since then, the threshold autoregressive (TAR) model has been applied widely across various domains. As detailed by Tong (1990), TAR models offer practical methodologies for exploring complex dynamical systems in time series, as they capture nonlinear dynamics, limit cycles, abrupt jumps, and asymmetries.
The primary goal of the TAR model is to identify threshold-driven regimes in time-varying systems and provide autoregressive representations of these regimes. The thresholding mechanism makes the TAR model particularly useful, as it decomposes complex stochastic systems into simpler subsystems, allowing for more efficient implementation and better interpretation. TAR models are well-suited to represent systems with shifting levels and sudden bursts, often seen in time series data such as EEG signals. For a comprehensive review of the applications of threshold nonlinearity and TAR models, we refer the reader to Hansen (2011) and Chen et al. (2011). In this study, we introduce the TAR model as an effective framework for analyzing the complex and highly dynamic brain activity captured through EEG signals.
For a stochastic process denoted as, $Y_{t}$, a general TAR model having $k$ regimes can be represented as follows:
@@
For example, if at time $t$, let $Z_{t-d}=a\in(r_{j-1},r_{j})$ then the active regime at that time is characterized by linear AR model;
@@
The overall process $Y_{t}$ is nonlinear if there are at least two regimes with different models, even if each is linear. Therefore, by describing the different regimes by means of threshold variables, we can ascertain the temporal behavior of the dynamics.
The unknown parameters for the model in Equation (1) can be assembled as follows:
@@
Since the TAR model is a piece-wise linear model, the estimates of the parameters can be readily obtained by using the least squares (LS) method conditional on hyperparameters (i.e., estimation is equivalent to the maximum likelihood estimation under the auxiliary assumption that $\varepsilon_{t}$ is i.i.d.$N(0,\sigma_{\varepsilon_{t}}^{2})$). Thus, the conditional least squares (CLS) method has a significant advantage in that it allows the exploration of the threshold parameter and hyperparameters without imposing any restrictions. Chan (1993) demonstrated the consistency of the CLS method under certain mild conditions, such as stationarity and the presence of a discontinuous conditional mean function.
@@
where $T$ denotes the sample size. The LS estimators $\widehat{\Omega}$ and the minimization of the objective function in Equation (2) can be achieved conditionally by a grid search considering all viable combinations of hyperparameters $\{j,p_{j},r,d|j=1,\ldots,k\}$. The optimization requires the execution of approximately $T^{k-1}/2\times d\times p_{(1)}\times\ldots\times p_{(k)}$ arranged autoregressions. Arranged autoregression is an autoregression where cases are rearranged by $\Gamma$-dependent sample splitting. Note that rearrangement of the sample based on the threshold values does not change the temporal structure of the series within cases (Tsay, 1989).
The asymptotic theory of LS estimation and inference for TAR models was addressed and developed by Chan (1993), Chan and Tsay (1998), Hansen (1997, 2000), Li and Ling (2012), Zhang et al. (2023).
@@ 2.3 Proposed Approach: TAR for causality in EEG networks
EEG functions as a network of interconnected components, where connectivity between node pairs can be the subject of investigation. This study proposes a 2-regime TAR model based approach to dissect causal interactions between EEG channels. The main objective is to map the net causal influence each node receives from others through pairwise connectivity analysis while minimizing interference from excluded nodes.
Let $N=\{Y^{*}_{t},X^{*}_{t},Z_{1,t},\ldots,Z_{n-2,t}\}$ represents $n$-dimensional network of multiple EEG time series; $Y^{*}$ and $X^{*}$ be the nodes of interest in this network, between which we aim to discover the causal relationship. Considering $Y$ and $X$ be the neutralized counterparts of $Y^{*}$ and $X^{*}$ from the impact of other nodes in the network, the ”net” activities of nodes $Y^{*}$ and $X^{*}$ at time $t$ can be denoted as $Y_{t}$ and $X_{t}$. Thereby, the 2-regime TAR model for the node $Y$, and the node $X$, which is treated as both the threshold and predictive variable, can be formulated as follows:
@@
where $u(X_{t-d})$ is the function constructs the threshold variable; $d$ denotes delay parameter with positive integer that establishes the time delay at which the threshold function $u(X_{t-d})$ significantly influences the structure of $Y_{t}$; $r\in[\text{min }u(X_{t-d}),\text{max }u(X_{t-d})]$ denotes threshold value; $p_{1}$ and $p_{2}$ indicates $1^{th}$ and $2^{nd}$ regime lag order of the node $Y_{t}$, respectively; $q_{1}$ and $q_{2}$ indicates $1^{th}$ and $2^{nd}$ regime lag order of the node $X_{t}$, respectively; $\{\phi_{0}^{(1)};\phi_{i}^{(1)}|i=1,2,\ldots,p_{1};\theta_{\ell}^{(1)}|\ell=1,2,\ldots,q_{1}\}$ are parameters for $1^{th}$ regime; and $\{\phi_{0}^{(2)};\phi_{i}^{(2)}|i=1,2,\ldots,p_{2};\theta_{\ell}^{(2)}|\ell=1,2,\ldots,q_{2}\}$ are parameters for $2^{nd}$ regime; $\{\varepsilon_{t}\}$ is a sequence of martingale differences satisfying ${\mathbb{E}}(\varepsilon_{t}|F_{t-1})=0$, ${sup}_{t}{\mathbb{E}}(|\varepsilon_{t}|^{\delta}|F_{t-1})<\infty$ $almost$ $surely$ for some $\delta>2$, where $F_{t-1}$ is the $\sigma$-field generated by $\{\varepsilon_{t-i}|i=1,2\ldots\}$. The unknown set of parameters for the model in Equation (4) can be shown as follows:
@@
The hyperparameters of the model can be collected as;
@@
The parameters of the model in (4) can be estimated by LS method conditional on hyperparameters similar to minimization of the objective function shown in (2) and (3).
In order to discover TAR model-based causative interactions in a network $N$ consisting of multiple sources of information, called nodes, it is helpful to specify the following definitions regarding (4) and (5).
@@
The first definition implies that the presence of a threshold non-linearity in the TAR model (4) would be sufficient enough to show that $X_{t}$ has a threshold conditional connectivity with $Y_{t}$. Note that, as per this definition, the set of parameters $\Theta$ is considered an empty set. In other words, even if the past of $X_{t}$ does not have a significant predictive effect, which would be provided by $\Theta$, the threshold functional of $X_{t-d}$ has a particular influence on the structure of $Y_{t}$ in the sense that it controls which regime $Y_{t}$ will follow. To explore the threshold-connectivity (TC) stated in definition 1, we utilize the threshold non-linearity test scheme suggested by Hansen (1996, 2000).
@@
Models given in (6) and (7), which is being the most restricted, have a nested structure. Thereby, comparing them using a model selection criterion is equivalent to a likelihood ratio test (Hansen, 1999).
Let $S_{\text{\tiny M1}}={\hat{\varepsilon}_{\text{\tiny M1}}}^{\prime}\times{\hat{\varepsilon}_{\text{\tiny M1}}}^{\text{}}$ and $S_{\text{\tiny M2}}=\hat{\varepsilon}^{\prime}_{\text{\tiny M2}}\times{\hat{\varepsilon}_{\text{\tiny M2}}}^{\text{}}$ are denoting sum of squared residuals obtained by LS estimation in (8) and (9), respectively; and let $r_{0}$ is denoting the true value of the threshold value $r$; then testing the existence of threshold non-linearity can be accomplished by the testing the hypothesis of $\text{H}_{0}:r=r_{0}$ considering the following likelihood ratio (LR) test statistic;
@@
(X threshold Granger causes Y if X has both threshold nonlinearity effect on Y and also having predictive power on Y)
With our second definition, we specify the threshold Granger causality (TGC) by incorporating the impact of the threshold node’s predictive ability in the TAR model. Once the threshold nonlinearity is verified, the form of the causal interactions between the two nodes of a network can be explored and represented by terms of the TAR model. In addition to the previously defined threshold-connectivity, in this second definition, the threshold node $X_{t-d}$, whose threshold connectivity to $Y_{t}$ is determined by the threshold non-linearity test, is included as a predictor of the current state of $Y_{t}$. Hence, the identification of TGC would require the presence of statistically significant coefficients for any of the time-lagged terms of the $X_{t}$. Given the possibility of multiple valid representations of causal interactions, it is necessary to establish a model selection criterion that avoids over-parameterization for a well-refined representation of the causal connections. To this end, we suggest employing the heteroscedasticity-robust Wald statistics provided by Hansen (1997) be used as a term to be penalized in model selection criteria. The Wald test for causality can be structured to test the predictive capacity of the threshold variable included as explanatory covariates in the model. It should be noted that Li (2006) was the first to mention the TGC nomenclature and made use of the Wald test, proposed and studied in detail by Hansen (1996, 1997) for the TAR model specification (i.e., testing the restricted vs unrestricted models). In his model, Li (2006) utilized the dependent variable itself as the threshold variable. Therefore, it is worth mentioning that whilst we adopt the common TGC nomenclature, our suggested methodology diverges in terms of the causal connectivity framework.
@@
where the alternative hypothesis discerns the TGC by stating that at least one coefficient is nonzero.
If the threshold value $r$ splitting the sample and delay $d$ are known, the pointwise heteroscedasticity-consistent Wald statistic can be shown as follows:
@@
$V(r,d)=\sum_{t=1}^{T}W_{t}(r,d)W_{t}(r,d)^{\prime}\hat{\varepsilon_{t}}^{2}$; $W_{t}(r,d)=\bigg{(}W_{t}^{\prime}I(u(X_{t-d})\leq r)\text{ }\text{ }W_{t}^{\prime}I(u(X_{t-d})>r)\bigg{)}^{\prime}$, then according Hansen (1996); Davies (1977, 1987) the appropriate testing of $\text{H}_{0}$ can be done by using following LR statistic
@@
with $K$ is the number of parameters in the model.
The model selection criterion based estimation of TGC can be defined as the maximization of the objective function given in (14),
@@
1
2Step 1. Neutralize interference from other nodes:
@@
      Any significant evidence against the hypothesis of linearity indicates ${\text{TC}}_{X\to Y}$. If the decision is in favor of threshold non-linearity, proceed to Step 3. Otherwise, or for different causality assessments, repeat Step 2 according to different threshold functionals.

### v2

@@ 2.1 Inference in high-dimensional networks of time series
Most phenomena in the real world are composed of many interconnected components. Statistical inference in such high-dimensional networks of time series continues to pose a significant analytical challenge. In such interconnected multi-node networks, if the primary purpose and scope of the research are to determine the connectivity between any specific two nodes, the confounding effects of the other nodes in the network cannot be ignored. Since the estimation and interpretation of multi-parameter models become more difficult as the number of parameters increases, instead of using a multi-parameter model, we adopt a practical and innovative dimensionality reduction approach proposed by Aslan and Ombao (2024) to determine the connectivity between two specific nodes accounting for the effect of all the other nodes in the network. This approach employs spectral dynamic principal component analysis (sDPCA) in the frequency domain to create a low-dimensional representation of all other nodes. This approach was originally developed in Brillinger (1974). The resulting principal component scores are subsequently removed from the nodes of interest, thus eliminating the confounding effects. This method ensures that the effects of other nodes are effectively isolated, allowing for a precise determination of connectivity between the two nodes.
Let $N=\{Y^{*}_{t},X^{*}_{t},Z_{1,t},\ldots,Z_{n-2,t}\}$ be a network of time series with $n$ nodes. In this paper, we consider a network representing brain activity through EEG recordings, where each component (i.e., node or channel) is a time series. Within this network $N$, we let $Y^{*}$, and $X^{*}$ represent any two nodes of interest. As demonstrated in Aslan and Ombao (2024), sDPCA in the frequency domain offers practical benefits and is highly effective for the alleviation of dimensionality problems in high-dimensional settings. Thereby, the interference-free counterparts of the nodes of interest can be obtained in the following manner:
@@
where $f_{z}(.)$ is a function that accounts for the dynamic variability of the excluded nodes. In practice, $f_{z}(.)$ serves as a filter and can take any form, provided it effectively summarizes the net discharge of the excluded nodes.
@@ 2.2 Background on the threshold autoregressive model
The concept of switching between regimes based on a change point was first introduced by Quandt (1958) in linear regression. Two decades later, Tong (1978) and Tong and Lim (1980) extended this idea to nonlinear time-series analysis through threshold non-linearity. Since then, the threshold autoregressive (TAR) model has been applied widely across various domains. TAR models offer practical methodologies for exploring complex dynamical systems in time series, as they capture nonlinear dynamics, limit cycles, abrupt jumps, and asymmetries. The primary purpose of the TAR model is to identify threshold-driven regimes in time-varying systems and provide autoregressive representations of these regimes. The thresholding mechanism makes the TAR model particularly useful, as it decomposes complex stochastic systems into simpler subsystems, allowing for more efficient implementation and better interpretation. TAR models are well-suited to represent systems with shifting levels and sudden bursts, often seen in time series data such as EEG signals. For a comprehensive review of the applications of threshold nonlinearity and TAR models, we refer the reader to Hansen (2011) and Chen et al. (2011). We introduce the TAR model as an effective analytical framework in this study to gain insight into the intricate and highly dynamic brain activity via EEG signals.
A stochastic process $Y_{t}$ is TAR with $k$ regimes if it can be represented as follows
@@
The overall process $Y_{t}$ is nonlinear if there are at least two regimes with different models, even if each is linear.
The unknown parameters for the model in Equation (1) can be assembled as follows
@@
Since the TAR model is a piece-wise linear model, the estimates of the parameters can be readily obtained by using the least squares (LS) method conditional on hyperparameters (i.e., estimation is equivalent to the maximum likelihood estimation under the auxiliary assumption that $\varepsilon_{t}$ is i.i.d.$N(0,\sigma_{\varepsilon_{t}}^{2})$). Thus, the conditional least squares (CLS) method has a significant advantage in that it allows the exploration of the threshold parameter and hyperparameters without imposing any restrictions. The consistency of the CLS method was demonstrated in Chan (1993) to hold under certain mild conditions such as stationarity and the presence of a discontinuous conditional mean function.
@@
where $T$ denotes the sample size. The LS estimators $\widehat{\Omega}$ and the minimization of the objective function in Equation (2) can be achieved conditionally by a grid search considering all viable combinations of hyperparameters $\{j,p_{j},r,d|j=1,\ldots,k\}$. The optimization requires the execution of approximately $T^{k-1}/2\times d\times p_{(1)}\times\ldots\times p_{(k)}$ arranged autoregressions. Arranged autoregression is an autoregression where cases are rearranged by $\Gamma$-dependent sample splitting. Note that rearrangement of the sample based on the threshold values does not change the temporal structure of the series within cases (Tsay, 1989). The asymptotic theory of LS estimation and inference for TAR models was addressed and developed by Chan (1993) and Zhang et al. (2023) (among many others).
@@ 2.3 Proposed Approach: TAR for causality in EEG networks
EEG functions as a network of interconnected components, where connectivity between node pairs can be the subject of investigation. This study proposes a 2-regime TAR model based approach to dissect causal interactions between EEG channels. The main objective is to map the net causal influence each node receives from others through pairwise connectivity analysis while minimizing interference from excluded nodes. Let $N=\{Y^{*}_{t},X^{*}_{t},Z_{1,t},\ldots,Z_{n-2,t}\}$ represents $n$-dimensional network of multiple EEG time series; $Y^{*}$ and $X^{*}$ be the nodes of interest in this network, between which we aim to discover the causal relationship. Considering $Y$ and $X$ be the neutralized counterparts of $Y^{*}$ and $X^{*}$ from the impact of other nodes in the network, the ”net” activities of nodes $Y^{*}$ and $X^{*}$ at time $t$ can be denoted as $Y_{t}$ and $X_{t}$. Thereby, the 2-regime TAR model for the node $Y$, and the node $X$, which is treated as both the threshold and predictive variable, can be formulated as follows:
@@
where $u(X_{t-d})$ is the function constructs the threshold variable; $d$ denotes delay parameter with positive integer that establishes the time delay at which the threshold function $u(X_{t-d})$ significantly influences the structure of $Y_{t}$; $r\in[\text{min }u(X_{t-d}),\text{max }u(X_{t-d})]$ denotes threshold value; $p_{1}$ and $p_{2}$ indicates $1^{th}$ and $2^{nd}$ regime lag order of the node $Y_{t}$, respectively; $q_{1}$ and $q_{2}$ indicates $1^{th}$ and $2^{nd}$ regime lag order of the node $X_{t}$, respectively; $\{\phi_{0}^{(1)};\phi_{i}^{(1)}|i=1,2,\ldots,p_{1};\theta_{\ell}^{(1)}|\ell=1,2,\ldots,q_{1}\}$ are parameters for $1^{th}$ regime; and $\{\phi_{0}^{(2)};\phi_{i}^{(2)}|i=1,2,\ldots,p_{2};\theta_{\ell}^{(2)}|\ell=1,2,\ldots,q_{2}\}$ are parameters for $2^{nd}$ regime; $\{\varepsilon_{t}\}$ is a sequence of martingale differences satisfying ${\mathbb{E}}(\varepsilon_{t}|F_{t-1})=0$, ${sup}_{t}{\mathbb{E}}(|\varepsilon_{t}|^{\delta}|F_{t-1})<\infty$ $almost$ $surely$ for some $\delta>2$, where $F_{t-1}$ is the $\sigma$-field generated by $\{\varepsilon_{t-i}|i=1,2\ldots\}$.
The unknown set of parameters for the model in Equation (4) can be shown as follows:
@@
The hyperparameters of the model can be collected as
@@
The parameters of the model in (4) can be estimated by LS method conditional on hyperparameters similar to minimization of the objective function shown in (2) and (3). To discover TAR model-based causative interactions in a network $N$ consisting of multiple sources of information, called nodes, it is helpful to specify the following definitions regarding (4) and (5).
@@
The first definition implies that the presence of a threshold non-linearity in the TAR model (4) would be sufficient to show that $X_{t}$ has a threshold conditional causal effect on $Y_{t}$. Note that, as per this definition, the set of parameters $\Theta$ is considered an empty set. In other words, even if the past of $X_{t}$ does not have a significant predictive effect, which would be provided by $\Theta$, the threshold functional of $X_{t-d}$ has a particular influence on the structure of $Y_{t}$ in the sense that it controls which regime $Y_{t}$ will follow. To explore the threshold-connectivity (TC) stated in definition 1, we utilize the threshold non-linearity test scheme suggested by Hansen (1996, 2000).
@@
Models given in (6) and (7), which are the most restricted, have a nested structure. Hence, comparing them using a model selection criterion is equivalent to a likelihood ratio test (Hansen, 1999).
Let $S_{\text{\tiny M1}}={\hat{\varepsilon}_{\text{\tiny M1}}}^{\prime}\times{\hat{\varepsilon}_{\text{\tiny M1}}}^{\text{}}$ and $S_{\text{\tiny M2}}=\hat{\varepsilon}^{\prime}_{\text{\tiny M2}}\times{\hat{\varepsilon}_{\text{\tiny M2}}}^{\text{}}$ denote the sum of squared residuals obtained by LS estimation in (8) and (9), respectively; and let $r_{0}$ is denoting the true value of the threshold value $r$; then testing the existence of threshold non-linearity can be accomplished by the testing the hypothesis of $\text{H}_{0}:r=r_{0}$ considering the following likelihood ratio (LR) test statistic;
@@
(X threshold Granger causes Y if X has both threshold nonlinearity effect on Y and also having predictive power on Y). The second definition specifies the threshold Granger causality (TGC) by incorporating the impact of the threshold node’s predictive ability in the model. Once the threshold nonlinearity is verified, the form of the causal interactions between the two nodes can be explored and represented in terms of the TAR model. In this second definition, the threshold node $X_{t-d}$, whose threshold connectivity to $Y_{t}$ is determined by the threshold non-linearity test, is included as a predictor of the current state of $Y_{t}$. Hence, the identification of TGC would require the presence of statistically significant coefficients for any of the time-lagged terms of the $X_{t}$. Given the possibility of multiple valid representations of causal interactions, it is necessary to establish a model selection criterion that avoids over-parameterization for a well-refined representation of the causal connections. To this end, we suggest employing the heteroscedasticity-robust Wald statistics provided by Hansen (1997) to be used as a term to be penalized in model selection criteria. The Wald test for causality can be structured to test the predictive capacity of the threshold variable included as explanatory covariates in the model. It should be noted that Li (2006) was the first to mention the TGC nomenclature and made use of the Wald test, proposed and studied in detail by Hansen (1996, 1997) for the TAR model specification (i.e., testing the restricted vs unrestricted models). In his model, Li (2006) utilized the dependent variable itself as the threshold variable. Therefore, it is worth mentioning that whilst we adopt the common TGC nomenclature, our suggested methodology diverges in terms of the causal connectivity framework.
@@
where the alternative hypothesis discerns the TGC by stating that at least one coefficient is nonzero. If the threshold value $r$ splitting the sample and delay $d$ are known, the pointwise heteroscedasticity-consistent Wald statistic can be shown as follows:
@@
$V(r,d)=\sum_{t=1}^{T}W_{t}(r,d)W_{t}(r,d)^{\prime}\hat{\varepsilon_{t}}^{2}$; $W_{t}(r,d)=\bigg{(}W_{t}^{\prime}I(u(X_{t-d})\leq r)\text{ }\text{ }W_{t}^{\prime}I(u(X_{t-d})>r)\bigg{)}^{\prime}$, then according Hansen (1996); Davies (1987) the appropriate testing of $\text{H}_{0}$ can be done by using following LR statistic
@@
with $K$ is the number of parameters in the model. The model selection criterion based estimation of TGC can be defined as the maximization of the objective function given in (14),
@@
3
4
  1. 5Step 1.
    Neutralize interference from other nodes:
@@
      Any significant evidence against the hypothesis of linearity indicates ${\text{TC}}_{X\to Y}$. If the decision is in favor of threshold non-linearity, proceed to the next step. Otherwise, or for different causality assessments, repeat Step 2 according to different threshold functionals.